# Generative Active Learning for Image Synthesis Personalization

### v1

@@ III-C Balancing the Exploitation and Exploration
As aforementioned, in the progression of GAL iterations, we need to keep the knowledge learned at past rounds while encouraging the model to explore. This introduces a exploitation-exploration dilemma [13]. To address this challenge, we propose evaluating the openness of the model at each round, using it as an indicator of the expected contribution of the novel information introduced in that round. Given that the novel information is encapsulated within the newly included reference images, we can utilize this indicator as a weight to regulate their impact on the learning process in the subsequent round. This encourages the exploration when the expected contribution is high, otherwise encourages the exploitation.
@@
To assess the openness of a round, we can utilize the uncertainty score previously computed using 6. Our rationale is that as the model explores more directions, its level of openness increases. Hence, the openness score for the round can be estimated by calculating
$\Delta(f_{\mathbf{\theta}})=\frac{\lambda}{|\mathcal{A}|}\sum_{i=1}^{|\mathcal{A}|}\Omega(\mathbf{a}_{i})$ (11)
@@
where $\lambda$ is a learning rate. This can be used to weight the newly include reference images to control their degrees of influence to the loss $L_{rec}$ (2).
@@
An additional outcome of 11 is its potential to establish an adaptive stopping criterion for GAL learning, in contrast to the fixed number of iterations often set in DAL. The concept behind this approach is to halt the learning process when there are fewer directions left to explore than anticipated. The stopping criteria is then simply written as
|$\Big{|}\{\Omega(\mathbf{a}_{i})\hskip 2.84544pt|\hskip 2.84544pt\Omega(\mathbf{a}_{i})>0,\mathbf{a}_{i}\in\mathcal{A}\}\Big{|}<k.$ (12)
TABLE I: The performance of different strategies.
@@
Uncertainty $+$ Balance (ours) 0.311 0.762 0.235 0.339 0.664 0.049
@@ IV Experiments
In style-driven personalization, we adopt the evaluation dataset used in the StyleDrop [10]. This dataset comprises various styles, such as watercolor painting, oil painting, 3D rendering, and cartoon illustration. 190 basic text prompts sourced from the Parti prompts dataset [40] are used to generate images, yielding a total of 36,480 images.
@@
In object-driven personalization, we adopt almost all concepts that have been previously used in related studies [27, 32], comprising a total of 10 categories including animals, furniture, containers, houses, plants, and toys. We use the 20 prompts in [27], which cover a wide range of test scenarios. In total, this process generates 6,400 images for a complete training cycle.
@@
Figure 2: Examples of images generated by anchor prompts with higher priority (left) and lower priority (right). Their features are highlighted in the feature space (middle). Poor-quality images that exhibit non-SoI are distributed near the reference, while high-quality images are located far from the reference. These images are from the anchor prompts in round 2.
@@
Evaluation Metrics. We utilize three metrics: 1) Text-alignment (TXT-ALN) assesses how well the generated images align with the intended textual descriptions This can be implemented by calculating the similarity between the CLIP image feature and the text feature. 2) Image-alignment (IMG-ALN) measures the extent to which the generated images capture the content or style present in the reference images. This can be implemented by the CLIP feature similarity between reference images and generated images. 3) Overfit (OVF) evaluates the proportion of overfitting in the test samples. Lower scores indicate better performance in terms of generalization and avoiding overfitting.
Base Model. In contrast to many closed-source ISP models developed by large companies, DreamBooth [9] is one of a few most widely used open-source methods with promising generation results. Therefore, we also employ DreamBooth as our base model. For our proposed method, we set the values of $m$, $k$, and $\lambda$ to 18, 3, and 0.0035, respectively.
IV-A Does generative active learning work in ISP?
@@
To evaluate the performance of different strategies on ISP, we compare our method with two commonly adopted querying strategies, including Random Sampling and Human Sampling. To be fair and efficient, we set a maximum number of rounds to 4 in all experiments. The initial round is based on original references without any synthetic data. The results are shown in I. It is evident from the results that both Random and Human strategies do not necessarily enhance the baseline performance. Instead, these strategies show a degradation on style-personalization of 22.3% (14.5%), 10.7%(10.4%), and 91.2%(24.0%) on TXT-ALN, IMG-ALN, and OVF, respectively. The unexpected degradation observed in the Human strategy, which is often considered an oracle in DAL, confirms the fundamental distinction between discriminative and generative tasks: while human annotators can easily differentiate between positive and negative samples, evaluating generalizability is a more challenging aspect. Therefore, we integrate human annotators with our balancing scheme to create a run that combines their selections and fairly weighs them for improved learning. The results are shown as the last row in I which demonstrates an approaching optimal performance and thus can be used as an oracle. Spuriously, our proposed method (Uncertainty+Balance) achieves a comparable performance with the oracle run. This has validated its effectiveness.
@@
Figure 3: Results of GAL over iterations. The images shown in the $1^{st}$ and $2^{nd}$ groups are for style- and object-driven ISP, respectively. The non-SoI and SoI are gradually disentangled and dragons or glasses are generated. Additional examples are available in the Appendix.
@@
Figure 4: The curves shown in the figure resemble clock arms extending from the baseline performance points. As these arms move in an anti-clockwise direction towards the top-right corners, better performance is observed.
IV-B How does the uncertainty sampling work?
@@
To gain deeper insights, we conduct a case study to observe the rationale behind the uncertainty sampling. 2 shows the distribution of images generated by the anchor prompts. Within the feature space, multiple sub-distributions can be observed. One particular distribution is centered around the reference, consisting of poor-quality images that exhibit non-SoI of the reference, such as the failure cases illustrated in 2. In contrast, images that align well are located far from the reference, forming smaller distributions that exclude non-SoI, like the successful samples of generated billboard, key, and candy images. Additionally, the distributions of good and bad samples across these three directions demonstrate significant diversity, suggesting a limited ability to generalize along these directions. As a result, these directions are given higher priority for querying based on our uncertainty metric. On the other hand, distributions at the directions of table and banana are homogeneous. Consequently, these directions exhibit lower entropy and lower querying priority. This observation aligns with the rationale we presented earlier.
IV-C How does GAL progress over iterations?
@@
To examine the progress of GAL over iterations, we present the performance in each round, as shown in 4, and visualize the evolution through the cases in 3. One notable observation is the dramatic and consistent decrease in performance of the Random strategy due to the inferior samples by random selection. After adopting a better querying strategy, the rate of decrease becomes much slower, and Uncertainty sampling begins to outperform the baseline on TXT-ALN for object-driven personalization, which suggests the effectiveness of valuable samples in enhancing generative models. The best overall progress is achieved through the combined strategies of Uncertainty sampling and the balancing scheme. We can find that TXT-ALN consistently improves and reaches its highest alignment in round 4, while IMG-ALN remains within a reasonable range. This trend is evident in 3, where the non-SoI semantics gradually disappear, and the number of successful generations of glasses placed on $S^{*}$ or dragon in style $S^{*}$ increases. Meanwhile, the SoI is maintained throughout the iteration rounds. These results indicate a progressive improvement by GAL as the iterations proceed. Additional comprehensive examples are available in the Appendix.
@@
TABLE II: Comparison with SOTA methods for object-driven ISP.
@@
+ UNC + BAL (R2) 0.306 0.777 0.270
+ UNC + BAL (R3) 0.305 0.782 0.293
+ UNC + BAL (R4) 0.311 0.762 0.235
+ ORACLE + BAL 0.307 0.772 0.254
TABLE III: Comparison with SOTA methods for style-driven ISP.
@@
+ UNC + BAL (R2) 0.328 0.679 0.096
+ UNC + BAL (R3) 0.337 0.666 0.056
+ UNC + BAL (R4) 0.339 0.664 0.049
+ ORACLE + BAL 0.342 0.650 0.023
@@
Figure 5: Qualitative comparison between our method and SOTA methods for personalized content generation. Our method produces text-aligned images compared with other methods. Besides, our method shows better preservation of details, including the color of cat statue and the spout of teapot. Additional comprehensive examples are available in the Appendix.
IV-D Comparison with SOTA methods
@@
For object driven-personalization, we compare 4 popular state-of-the-art (SOTA) methods including Textual Inversion [32], Custom Diffusion [27], DreamBooth [9], IP-Adapter [25]. The results are shown in II. Compared to IP-Adapter, Textual Inversion, and Custom Diffusion, our method demonstrates significant improvements on TXT-ALN and OVF throughout all rounds, achieving 15.2%(68.0%), 12.3%(46.7%), and 3.3%(18.1%) on TXT-ALN (OVF) in terms of round 4. Since the non-SoI semantics dominate the outputs of the other approaches, our method exhibits a slight decrease on IMG-ALN. 5 provides visual evidence of our method’s superior text and object fidelity. The success in higher text fidelity can be observed in the accurate placement of the cat statue in the Grand Canyon and the realistic interaction between the marigold flowers and the teapot. Furthermore, our method enhances object fidelity by accurately reconstructing only one spout and better preserving the color of the cat statue.
@@
For style-driven personalization, we conduct a comparison between four variations of StyleDrop [10]: base model, random feedback (StyleDrop-Random), clip-based feedback (StyleDrop-CF), and human feedback (StyleDrop-HF). Additionally, we include the results of DreamBooth on Imagen [23] as well as other pre-trained models like Imagen and Muse [41], as reported by [10]. It is clear that our method significantly outperforms the pre-trained models and achieves superior performance in terms of 5.3% and 3.0% on TXT-ALN compared to the dedicated human feedback and clip-based feedback of StyleDrop. It is worth noting that the closed-source StyleDrop is built on a more powerful backbone, Muse, compared to Stable Diffusion. This indicates that the open-sourced ISP models are able to achieve better performance with GAL.
V Conclusion

### v2

@@ 3.3. Balancing the Exploitation and Exploration
As aforementioned, in the progression of GAL iterations, we need to keep the knowledge learned at past rounds while encouraging the model to explore. This introduces an exploitation-exploration dilemma (Wei and Yang, 2011). To address this challenge, we propose evaluating the openness of the model at each round, using it as an indicator of the expected contribution of the novel information introduced in that round. Given that the novel information is encapsulated within the newly included reference images, we can utilize this indicator as a weight to regulate their impact on the learning process in the subsequent round. This encourages the exploration when the expected contribution is high, otherwise encourages the exploitation.
@@
To assess the openness of a round, we can utilize the uncertainty score previously computed by Equation 6. Our rationale is that as the model explores more directions, its level of openness increases. Hence, the openness score for the round can be estimated by calculating
(11) $\Delta(f_{\mathbf{\theta}})=\frac{\lambda}{|\mathcal{A}|}\sum_{i=1}^{|\mathcal{A}|}\Omega(\mathbf{a}_{i})$
@@
where $\lambda$ is a learning rate. This can be used to weight the newly include reference images to control their degrees of influence to the loss $L_{rec}$ (Equation 2).
@@
An additional outcome of Equation 11 is its potential to establish an adaptive stopping criterion for GAL learning, in contrast to the fixed number of iterations often set in DAL. The concept behind this approach is to halt the learning process when there are fewer directions left to explore than anticipated. The stopping criteria is then simply written as
(12) |$\Big{|}\{\Omega(\mathbf{a}_{i})\hskip 2.84544pt|\hskip 2.84544pt\Omega(\mathbf{a}_{i})>0,\mathbf{a}_{i}\in\mathcal{A}\}\Big{|}<k.$
Table 1. The performance of different GAL strategies including random selection (Random), human feedback (Human), direction-based uncertainty sampling (Uncertainty), direction-based uncertainty sampling with balance scheme (Uncertainty $+$ Balance), human feedback with balance scheme (Human $+$ Balance).
@@
Uncertainty $+$ Balance (ours) 0.309 0.771 0.268 0.337 0.669 0.058
@@ 4. Experiments
For style-driven ISP, we adopt the evaluation dataset used in the StyleDrop (Sohn et al., 2023). This dataset comprises various styles, such as watercolor painting, oil painting, 3D rendering, and cartoon illustration. 190 basic text prompts sourced from the Parti prompts dataset (Yu et al., 2022) are used to generate images, yielding 36,480 images.
@@
For object-driven ISP, we adopt almost all concepts that have been previously used in related studies (Kumari et al., 2023; Gal et al., 2022), comprising a total of 10 categories including animals, furniture, containers, houses, plants, and toys. We use the 20 prompts in (Kumari et al., 2023), which cover a wide range of test scenarios. In total, this process generates 6,400 images for a complete training cycle.
@@
Figure 3. Examples of images generated by anchor prompts in round 2 with higher priority (left) and lower priority (right). Their CLIP image features are highlighted in the tSNE (Van der Maaten and Hinton, 2008) space (middle). Poor-quality images that exhibit non-SoI are distributed near the reference, while high-quality images are located far from the reference.
@@
Evaluation Metrics. We utilize three metrics: 1) Text-alignment (TXT-ALN) assesses how well the generated images align with the intended textual descriptions This can be implemented by calculating the similarity between the CLIP image feature and the text feature. 2) Image-alignment (IMG-ALN) measures the extent to which the generated images capture the content or style present in the reference images. This can be implemented by the CLIP feature similarity between reference images and generated images. 3) Overfit (OVF) evaluates the proportion of overfitting in the test samples based on Equation 10. Lower scores indicate better performance in terms of generalization and avoiding overfitting.
Base Model. DreamBooth (Ruiz et al., 2023) is a widely adopted method with promising generation results. Thus, we utilize DreamBooth as our baseline, with the first-round results derived directly from it without synthetic training data. For our proposed method, we set the values of $m$ and $\lambda$ to 10 and 0.005, respectively. The initial anchor directions comprise 18 prompts. We select top-3, along with their associated highest-fidelity images, to serve as additional training pairs. We provide more implementation details in the Appendix.
4.1. Does generative active learning work in ISP?
@@
To evaluate the performance of different strategies, we compare our method with two commonly adopted querying strategies, including Random Sampling and Human Sampling. To be fair and efficient, we set a maximum number of rounds to 4 in all experiments. The initial round is based on original references without any synthetic data. The results are shown in Table 1. It is evident from the results that both Random and Human strategies do not necessarily enhance the baseline performance. Instead, these strategies show a degradation on style-personalization of 22.3% (14.5%), 10.7% (10.4%), and 91.2% (24.0%) on TXT-ALN, IMG-ALN, and OVF, respectively. The unexpected degradation observed in the Human strategy, which is often considered an oracle in DAL, confirms the fundamental distinction between discriminative and generative tasks: while human annotators can easily differentiate between positive and negative samples, evaluating generalizability is a more challenging aspect. Therefore, we integrate human annotators with our balancing scheme to create a run that combines their selections and fairly weighs them for improved learning. The results are shown as the last row in Table 1 which demonstrates an approaching optimal performance and thus can be used as an oracle. Spuriously, our proposed method (Uncertainty+Balance) achieves a comparable performance with the oracle run. This validates its effectiveness.
@@
Figure 4. Results of GAL over iterations. The images shown in the $1^{st}$ and $2^{nd}$ groups are for style- and object-driven ISP, respectively. The non-SoI and SoI are gradually disentangled and dragons or glasses are generated. Additional examples are available in the Appendix.
@@
Figure 5. The curves shown in the figure resemble clock arms extending from the baseline performance points. As these arms move in an anti-clockwise direction towards the top-right corners, better performance is observed.
4.2. How does the uncertainty sampling work?
@@
To gain deeper insights, we conduct a case study to observe the rationale behind the uncertainty sampling. Figure 3 shows the distribution of images generated by the anchor prompts. Within the feature space, multiple sub-distributions can be observed. One particular distribution is centered around the reference, consisting of poor-quality images that exhibit non-SoI of the reference, such as the failure cases illustrated in Figure 3. In contrast, images that align well are located far from the reference, forming smaller distributions that exclude non-SoI, like the successful samples of generated billboard, key, and candy images. Additionally, the distributions of good and bad samples across these three directions demonstrate significant diversity, suggesting a limited ability to generalize along these directions. As a result, these directions are given higher priority for querying based on our uncertainty metric. On the other hand, distributions at the directions of table and banana are homogeneous. Consequently, these directions exhibit lower entropy and lower querying priority. This observation aligns with the rationale we presented earlier.
4.3. How does GAL progress over iterations?
@@
To examine the progress of GAL over iterations, we present the performance in each round, as shown in Figure 5, and visualize the evolution through the cases in Figure 4. One notable observation is the dramatic and consistent decrease in performance of the Random strategy due to the inferior samples by random selection. After adopting a better querying strategy, the rate of decrease becomes much slower, and Uncertainty sampling begins to outperform the baseline on TXT-ALN for object-driven personalization, which suggests the effectiveness of valuable samples in enhancing generative models. The best overall progress is achieved through the combined strategies of Uncertainty sampling and the balancing scheme. We can find that TXT-ALN consistently improves and reaches its highest alignment in round 4, while IMG-ALN remains within a reasonable range. This trend is evident in Figure 4, where the non-SoI semantics gradually disappear, and the number of successful generations of glasses placed on $S^{*}$ or dragon in style $S^{*}$ increases. Meanwhile, the SoI is maintained throughout the iteration rounds. These results indicate a progressive improvement by GAL as the iterations proceed. Additional comprehensive examples are available in the Appendix.
@@
Table 2. Comparison with SOTA methods for object-driven ISP. Results marked with indicate our re-implementation using publicly available codebases.
@@
$+$ Uncertainty $+$ Balance (R2) 0.304 0.781 0.300
$+$ Uncertainty $+$ Balance (R3) 0.308 0.769 0.248
$+$ Uncertainty $+$ Balance (R4) 0.309 0.771 0.268
$+$ Oracle $+$ Bablance (R4) 0.307 0.772 0.254
Table 3. Comparison with SOTA methods for style-driven ISP. Results marked with are obtained from (Sohn et al., 2023).
@@
$+$ Uncertainty $+$ Balance (R2) 0.328 0.683 0.097
$+$ Uncertainty $+$ Balance (R3) 0.336 0.671 0.059
$+$ Uncertainty $+$ Balance (R4) 0.337 0.669 0.058
$+$ Oracle $+$ Balance (R4) 0.342 0.650 0.023
@@
Figure 6. Qualitative comparison between our method and SOTA methods for personalized content generation. Our method produces text-aligned images compared with other methods. Additional comprehensive examples are available in the Appendix.
4.4. Comparison with SOTA methods
@@
For object driven-personalization, we compare 4 popular state-of-the-art (SOTA) methods including Textual Inversion (Gal et al., 2022), Custom Diffusion (Kumari et al., 2023), DreamBooth (Ruiz et al., 2023), IP-Adapter (Ye et al., 2023). The results are shown in Table 2. Compared to IP-Adapter, Textual Inversion, and Custom Diffusion, our method demonstrates significant improvements on TXT-ALN and OVF throughout almost all rounds, achieving 14.4%(63.5%), 11.6%(39.2%), and 2.7%(6.6%) on TXT-ALN (OVF) in terms of round 4. Since the non-SoI semantics dominate the outputs of the other approaches, our method exhibits a slight decrease on IMG-ALN. Figure 6 provides visual evidence of our method’s superior text and object fidelity. The success in higher text fidelity can be observed in the accurate placement of the cat statue in the Grand Canyon and the realistic interaction between the marigold flowers and the teapot. Furthermore, our method enhances object fidelity by accurately reconstructing only one spout and better preserving the color of the cat statue.
@@
For style-driven personalization, we conduct a comparison between four variations of StyleDrop (Sohn et al., 2023): base model, random feedback (StyleDrop-Random), clip-based feedback (StyleDrop-CF), and human feedback (StyleDrop-HF). Additionally, we include the results of DreamBooth on Imagen (Saharia et al., 2022) as well as other pre-trained models like Imagen and Muse (Chang et al., 2023), as reported by (Sohn et al., 2023). It is clear that our method significantly outperforms the pre-trained models and achieves superior performance in terms of 4.7% and 2.4% on TXT-ALN compared to the dedicated human feedback and clip-based feedback of StyleDrop. It is worth noting that the closed-source StyleDrop is built on a more powerful backbone, Muse, compared to Stable Diffusion. This indicates that the open-sourced ISP models are able to achieve better performance with GAL.
Table 4. The percentage of user preference on our proposed method (Uncertainty + Balance) compared to Round 1 (DreamBooth) and Oracle feedback (Human + Balance).
Object Style
TXT-ALN IMG-ALN TXT-ALN IMG-ALN
Ours vs. Round 1 60.4 % 32.5 % 77.8% 59.8%
Ours vs. Oracle 53.8 % 46.2 % 47.0% 67.8%
4.5. User Study
We conduct a user study involving two comparison tasks. Participants are presented with reference images and a text prompt, and are asked to choose the more faithful result in terms of object/style and text fidelity. This process yields a total of 4800 responses from 8 participants. The results are shown in Table 4. It is clear that our method significantly improves the text alignment, with particularly notable gains in style-driven ISP where both text and style fidelity surpass round 1. This indicates the superior performance of GAL when users can only provide fewer samples. By comparing the oracle feedback, our automatic uncertainty sampling strategy performs comparable results. Notably, a majority of users prefer our style renderings rather than those trained from human selection. This further validates the effectiveness of our method. More details are available in Appendix.
Figure 7. Illustration of ablation experiments on style-driven ISP. (a) Variation in performance with the parameter $\lambda$. (b) Effects of different anchor set sizes. (c) Impact of selecting the top-$k$ prompts per iteration. (d) Results from varying the prompt composition within the anchor set.
4.6. Ablations
Figure 7 presents ablation studies on style-driven ISP to evaluate our model’s sensitivity to various hyperparameters. Details on the ablations for object-driven ISP are provided in the Appendix.
Learning Rate $\lambda$ on Openness. Subfigures (a) depicts the effect of the learning rate $\lambda$, which controls the scale of the openness score in Equation 11. It is obvious that a relatively higher $\lambda$ does not exhibit promising results, particularly in scenarios with a single reference for style-driven ISP. Additionally, a $\lambda$ below 0.05 results in stable performance.
Size of Anchor Set. As illustrated in subfigures (b), increasing the size of the anchor embedding set enhances style fidelity but reduces text alignment. Conversely, a smaller anchor size exhibits the opposite effect. Therefore, we consider a moderate size of 18 as our default setting.
Top-$k$ Anchor Prompts. Because of the trade-off between IMG-ALN and TXT-ALN metrics, as shown in subfigures (c), there is no globally optimal top-$k$ setting. Consequently, we adopt the top-3 selection as a standard practice based on relative performance.
Anchor Set Variability. Finally, we change the prompts in the anchor set, forming 3 distinct sets, each differing by at least 50%. As shown in subfigures (d), the results reveal our model’s robustness against variations in anchor prompts. This indicates the effectiveness of our uncertainty sampling method which selects the most constructive direction for model training.
5. Conclusion
@@
Appendix A Additional Implementation Details
For a fair comparison, we utilize the third-party implementation of HuggingFace for all SOTA methods in our experiments. We use Stable Diffusion 1.5 version as the pre-trained model. Besides, we apply 50 steps of DDPM (Denoising Diffusion Probabilistic Model) sampling with a guidance scale of 7.5 in all experiments. All experiments are conducted using A-100 GPUs.
In each round, we adopt DreamBooth to retrain the diffusion model. Specifically, the parameters of the text Transformer are frozen and the U-Net diffusion model is fine-tuned with a learning rate $5\times 10^{-6}$ and the batch size is 1. The training step for content learning is 800 steps (wooden pot for 600 steps to avoid overfitting) and style is 500 steps.
Appendix B Text Prompts for Additional Training Samples Generation
Here we provide the prompts used to synthesize the images for iterative training in Table 5 and Table 6. In each round, 3 prompts paired images will be selected as additional training samples.
Table 5. The anchor prompts of object personalization to generate synthetic samples.
$S^{*}$ in park” $S^{*}$ near a pool” $S^{*}$ on street”
$S^{*}$ in a forest” $S^{*}$ in a kitchen” $S^{*}$ on the beach”
$S^{*}$ in a restaurant” $S^{*}$ on snowy ground” $S^{*}$ on the moon”
$S^{*}$ in desert” $S^{*}$ in library” $S^{*}$ under the night sky”
$S^{*}$ in front of the castle” $S^{*}$ on the bridge” $S^{*}$ next to waterfall”
$S^{*}$ in the cave” $S^{*}$ at a concert venue” $S^{*}$ on a balcony”
Table 6. The anchor prompts of style personalization to generate synthetic samples.
“A tie in style $S^{*}$ “A table in style $S^{*}$ “A broccoli in style $S^{*}$
“A box in style $S^{*}$ “Candy in style $S^{*}$ “A billboard in style $S^{*}$
“A drum in style $S^{*}$ “A cell phone in style $S^{*}$ “An empty bottle in style $S^{*}$
“A glass in style $S^{*}$ “A golden key in style $S^{*}$ “A koala in style $S^{*}$
“A toy in style $S^{*}$ “A skateboard in style $S^{*}$ “An apple on the table in style $S^{*}$
“A bridge in style $S^{*}$ “Beach scene in style $S^{*}$ “A banana on the table in style $S^{*}$
Appendix C Stopping Criterion
Figure 8 illustrates a case study to show the effectiveness of the early stopping setting. We can find that the personalized model performs well toward text fidelity and image fidelity in the termination round. However, beyond this point, introducing non-informative samples results in a decline in performance. We can find the prominent edge features and unnatural brushstrokes in the generated outputs. These deteriorated results suggest that the model struggles to maintain its initial level of quality when confronted with less informative prompts. The stopping criterion plays an important role in preventing the inclusion of samples that could negatively impact its output quality when there are few anchor directions to explore.
Figure 8. Comparison between the termination round and extended round in the iterative training process. The details are better preserved in the termination round.
Figure 9. Illustration of ablation experiments on object-driven ISP. (a) Variation in performance with the parameter $\lambda$. (b) Effects of different anchor set sizes. (c) Impact of selecting the top-$k$ prompts per iteration. (d) Results from varying the prompt composition within the anchor set.
Appendix D Ablations on object-driven ISP
We provide ablations on object-driven ISP. As shown in Figure 9, we can draw similar conclusions on these hyperparameter selections.
Appendix E Details on User study
In the user study conducted in Section 4.5, we perform 2 pair comparison tasks between Final Round (Ours) and Round 1, Final Round (Ours) and Oracle (Human $+$ Balance). For each of 10 Objects and 5 Style, 20 queries are presented, which results in a total of $2\times(10\times 20+5\times 20)$ = 600 queries. In each query, 2 questions are included for assessing the object/style and text fidelity, respectively. Totally, 4800 responses are collected from 8 participants.
The user study interface is shown in Figure 10. It includes the reference object/style image, prompt text, 2 image candidates, 2 questions with 3 options for each, instruction on tasks, and other functional components. The order of the queries and image candidates is randomly set for unbiased comparisons.
Questions:
  • Task1. Choose one image that most aligns with the style of / object in the reference image.
    • Options: Image A, Equal, Image B
  • Task2. Choose one image that most describes the Prompt text.
    • Options: Image A, Equal, Image B
Instructions:
  • Task 1: First view the reference object/style image, then choose one image (Image A, or Image B) that most aligns with the style of / object in the reference image. If you cannot determine, choose Equal.
  • Task 2: Then view the Prompt text. Choose one image (Image A, or Image B) that most describes the Prompt text. If you cannot determine, choose Equal.
Appendix F Additional Results
This section provides samples of generated images. Figure 11 and Figure 12 show the generated images in each round based on object and style references. Figure 13 gives a more visualized comparison with SOTA methods. Figure 14, Figure 15, and Figure 16 show the generalization of our method on a wide range of prompts.
Figure 10. The interface for user preference study. The references and a generation prompt are shown at the top. Participants are required to select their preferred output by following the given instructions.
Figure 11. Examples of images generated in each round. The model in round 1 fails to produce the desired subjects, including food, sword, butterfly, and car. These subjects are successfully generated after using generative active learning.
Figure 12. Examples of images generated in each round. The model in round 1 fails to produce the desired subjects, including chocolate cake, marigold flowers, crochet, and orange sofa. These subjects are successfully generated after using generative active learning.
Figure 13. Comparison with other SOTA methods. Our method achieves better text and object fidelity.
Figure 14. Generated images using style references. Our method generalizes well on a diverse range of prompts.
Figure 15. Generated images using object references. Our method can be used for a variety of personalization purposes, such as background replacement, artistic rendering, and attribute editing.
Figure 16. Generated images using object references. Our method can be used for a variety of personalization purposes, such as background replacement, artistic rendering, and attribute editing.